OpenStudy (anonymous):

how many local maxima and minima does the polynomial y=-5x^2 +6x+9 have? maxima______ minima______

OpenStudy (anonymous):

no min max at \[x=\frac{-b}{2a}\]

OpenStudy (anonymous):

answer to "how many" is one and none respectively

OpenStudy (anonymous):

differentiate function. dy/dx=-10x + 6. differentiate one more time to know whether it's max/min. You get d^2y/dx^2 = -10, which is negative, so it's a maximum

OpenStudy (anonymous):

by the way all functions of the form ax^2+bx+c always have only 1 maxima/minima; if a is negative it's a maxima, if a is positive it's a minima.

OpenStudy (anonymous):

how did u find that sattelite

OpenStudy (anonymous):

If u diff ax^2 +bx + c -> 2ax +b = 0 for min/max ie x= -b/2a That is where the formula comes from.